Prove that y(x) = C1 sin2x + C2 cos2x is a solution of y


(2) + 4y = 0

1
Expert's answer
2021-11-15T17:56:16-0500

"\text{Given } y(x) = c_1 \sin 2x + c_2 \cos 2x \\
y'(x) = 2c_1 \cos 2x - 2c_2 \sin 2x \\
y''(x)= -4c_1 \cos 2x - 4c_2 \cos 2x \\
\implies y''(x) + 4y(x) = -4c_1 \cos 2x - 4c_2 \cos 2x + 4c_1 \sin 2x + 4c_2 \cos 2x \\
\qquad \qquad \qquad \qquad \quad = -4c_1 \cos 2x + 4c_1 \cos 2x -4c_1 \sin 2x + 4c_1 \cos 2x\\ 
 \qquad \qquad \qquad \qquad \quad = 0 \\
\therefore y(x) = c_1 \sin 2x + c_2 \cos 2x \text{ is a solution of } y''(x) + 4y(x) = 0"